OpenStudy (anonymous):

What is the value of f[g(5)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (anonymous):

can you find g(5) ?

OpenStudy (anonymous):

\[\LARGE g(x)=2x-5\] \[\LARGE g(5)=2\cdot 5-5\] ... so g(5)=?

OpenStudy (anonymous):

5